I do not understand this problem ???
Use elimination to solve:
x/2+2y=9
2x-y=18
please help.

Answer by longjonsilver(2297)

multiply top equation by 4 to give 2x + 8y = 36.

The 2 equations are therefore:
2x + 8y = 36
2x - y = 18

Subtract them, leaving us with 9y = 18. So y = 18/9 --> y = 2

Put this into either equation:
2x-y = 18
2x-(2) = 18
2x-2 = 18
2x = 20
x = 20/2
x = 10

Put both in the other equation to check your answer is correct:
x/2 + 2y = 9
10/2 + 2(2)
5 + 4
9 --> correct!

jon.